A buffer system is prepared by mixing 2 mols of HF(aq) (Ka=7.1 x 10^-4) and 1.8 moles NaF(aq) in a 500 mL solution. A) calculate the pH of the buffer. B) calculate the pH of the buffer after adding 0.2 mols of HCl. C) calculate the pH of the buffer after adding 0.3 mols LiOH


1
Expert's answer
2022-06-13T01:05:22-0400

CM (HF) = 2mol / 0,5 liters = 4 mol/l

CM (NaF) = 1,8mol / 0,5 liters = 3,6 mol/l

Ka = 7,1*10-4


A.)

pKa = -log (Ka)

pKa = -log(7,1*10-4) = 3,15


Henderson-Hasselbalch equation:

"pH= pK_a + \cfrac {[A^-]}{[HA]}"


"pH = 3,15 + \cfrac {3,6}{4}= \bold{4,05}"


B.)

CM(HCl) = 0,2 mol / 0,5 litres = 0,4 mol/l


NaF + HCl = NaCl + HF

3,6-0,4 – 0,4-0,4 – 0,4 – 4 + 0,4

[A] = 3,2 mol/l

[HA] = 4,4 mol/l


"pH= 3,15+ \cfrac {3,2}{4,4}= \bold{3,88}"


C.)

CM(LiOH) = 0,3mol / 0,5 litres = 0,6 mol/l


HF + LiOH = LiF + H2O

4-0,6 – 0,6-0,6 – 0,6 – 0,6


[A] = 3,6 + 0,6 = 4,2

[HA] = 4 - 0,6 = 3,4


"pH= 3,15+ \cfrac {4,2}{3,4}= \bold{4,385}"